OpenStudy (anonymous):

When I tried to solve the system 2x + y = 9 3x + 4y = 26 I ended up getting x = 1.3, I'm not sure that correct and I need help working through this problem!

OpenStudy (anonymous):

You should be getting x=2

OpenStudy (johnweldon1993):

Hmm I'm gonna use the elimination method here to solve...multiply the top equation by 4 and subtract the 2 equations 8x + 4y = 36 -3x + 4y = 26 -------------- 5x = 10 x = 2 is what you should be getting

OpenStudy (anonymous):

shouldn't that be a negative 4?

OpenStudy (johnweldon1993):

In my equation @hereweg0 ? I think I know what your saying...that is just a subtraction sign...not a negative sign

OpenStudy (johnweldon1993):

8x + 4y = 36 -(3x + 4y = 26) ---------------- 5x = 10 There...better to clear it up...

OpenStudy (anonymous):

okay yeah I understand. it would make sense if there was ( )

OpenStudy (anonymous):

Okay, I see my mistake! Thanks so much! (:

OpenStudy (johnweldon1993):

No problem! :)

OpenStudy (texaschic101):

you could also use substitution on this one, if you prefer

OpenStudy (anonymous):

@texaschic101 yeah I did that the first time, but I think elimination was the easier method